# Mixtures of Experts Unlock Parameter Scaling for Deep RL

### v2

@@ 2.1 Reinforcement Learning
$V^{\pi}(x):=\underset{\pi,\mathcal{P}}{\mathbb{E}}\left[\sum_{t=i}^{\infty}\gamma^{t}\mathcal{R}\left(x_{t},a_{t}\right)\mid x_{0}=x\right].$
The state-action function $Q^{\pi}$ quantifies the value of first taking action $a$ from state $x$, and then following $\pi$ thereafter: $Q^{\pi}(x,a):=\mathcal{R}(x,a)+\gamma\underset{x^{\prime}\sim\mathcal{P}(x,a)}{\mathbb{E}}V^{\pi}(x^{\prime})$. Every MDP admits an optimal policy $\pi^{*}$ in the sense that $V^{\pi^{*}}:=V^{*}\geq V^{\pi}$ uniformly over $\mathcal{X}$ for all policies $\pi$.
@@ 4 Empirical evaluation
Our implementation, included with this submission, is built on the Dopamine library (Castro et al., 2018) [^3] Dopamine code available at https://github.com/google/dopamine., which adds stochasticity (via sticky actions) to the ALE (Machado et al., 2018). We use the recommendations from Agarwal et al. (2021) for statistically-robust performance evaluations, in particular focusing on interquartile mean (IQM). Every experiment was run for 200M environment steps, unless reported otherwise, with 5 independent seeds, and we report 95% stratified bootstrap confidence intervals. All experiments were run on NVIDIA Tesla P100 GPUs, and each took on average 4 days to complete.
@@ 4.1 Soft MoE Helps Parameter Scalability
As Figure 1 demonstrates, Soft MoE provides clear performance gains, and these gains increase with the number of experts; for instance in Rainbow, increasing the number of experts from 1 to 8 results in a 20% performance improvement. In contrast, the performance of the base architectures declines as we widen its layer; for instance in Rainbow, as we increase the layer multiplier from 1 to 8 there is a performance decrease of around 40%. This is a finding consistent with prior work demonstrating the difficulty in scaling up deep RL networks (Farebrother et al., 2022; Taiga et al., 2022; Schwarzer et al., 2023). Top1-MoE seems to provide gains when incorporated into Rainbow, but fails to exhibit the parameter-scalability we observe with Soft MoE.
@@ 5.1 Offline RL
We begin by incorporating MoEs in offline RL, where agents are trained on a fixed dataset without environment interactions. Following prior work (Kumar et al., 2021b), we train the agents over 17 games and 5 seeds for 200 iterations, where 1 iteration corresponds to 62,500 gradient updates. We evaluated on datasets composed of 5%, 10%, 50% of the samples (drawn randomly) from the set of all environment interactions collected by a DQN agent trained for 200M steps (Agarwal et al., 2020). In Figure 11 we observe that combining Soft MoE with two modern offline RL algorithms (CQL (Kumar et al., 2020) and CQL+C51 (Kumar et al., 2022)) attains the best aggregate final performance. Top1-MoE provides performance improvements when used in conjunction with CQL+C51, but not with CQL. Similar improvements can be observed when using 10% of the samples (see Figure 17).
@@ 6 Related Work
Another line of research demonstrates that RL networks can be trained with a high sparsity level ($\sim$90%) without loss in performance (Tan et al., 2022; Sokar et al., 2022; Graesser et al., 2022; Obando-Ceron et al., 2024). These observations call for techniques to better utilize the network parameters in RL training, such as using MoEs, which we show decreases dormant neurons drastically over multiple tasks and architectures. To enable scaling networks in deep RL, prior works focus on algorithmic methods (Farebrother et al., 2022; Taiga et al., 2022; Schwarzer et al., 2023; Farebrother et al., 2024). In contrast, we focus on alternative network topologies to enable robustness towards scaling.
@@ 7 Discussion and Conclusion
Figure 12: Normalized performance across 26 Atari games for DrQ($\epsilon$) (left) and DER (right), with the ResNet architecture (Espeholt et al., 2018) and 8 experts (see Figure 16, for 4 experts). Soft MoE not only remains generally stable with more training, but also attains higher final performance. We report interquantile mean performance with error bars indicating 95% confidence intervals.
@@ Acknowledgements
The authors would like to thank Gheorghe Comanici, Owen He, Alex Muzio, Adrien Ali Taïga, Rishabh Agarwal, Hugo Larochelle, Ayoub Echchahed, and the rest of the Google DeepMind Montreal team for valuable discussions during the preparation of this work; Gheorghe deserves a special mention for providing us valuable feedback on an early draft of the paper. We thank the anonymous reviewers for their valuable help in improving our manuscript. We would also like to thank the Python community (Van Rossum & Drake Jr, 1995; Oliphant, 2007) for developing tools that enabled this work, including NumPy (Harris et al., 2020), Matplotlib (Hunter, 2007) and JAX (Bradbury et al., 2018).
@@ Appendix A Experimental details
Unless otherwise specified, in all experiments below we report the interquantile mean after 40 million environment steps; error bars indicate 95% stratified bootstrap confidence intervals (Agarwal et al., 2021). Most of our experiments were run with 20 games from the ALE suite (Bellemare et al., 2013a), as suggested by Fedus et al. (2020). However, for the Atari 100k agents (subsection 5.2), we used the standard set of 26 games (Kaiser et al., 2020) to be consistent with the benchmark. Finally, we also ran some experiments with the full set of 60 games. The specific games are detailed below.
@@ Appendix B Hyper-parameters list
Default hyper-parameter settings for DER (Van Hasselt et al., 2019) and DrQ($\epsilon$) (Kaiser et al., 2020; Agarwal et al., 2021). Table 1 shows the default values for each hyper-parameter across all the Atari games.
@@
Exploration $\epsilon$ 0.00015 0.00015
Learning Rate 0.01 0.01
@@
Default hyper-parameter settings for DQN (Mnih et al., 2015), Rainbow (Hessel et al., 2018). Table 2 shows the default values for each hyper-parameter across all the Atari games.
Table 2: Default hyper-parameters setting for DQN, Rainbow, IQN, Munchausen-IQN agents.
@@
Hyper-parameter DQN Rainbow IQN M-IQN
Adam’s($\epsilon$) 1.5e-4 1.5e-4 1.5e-4 1.5e-4
Batch Size 32 32 32 32
Conv. Activation Function ReLU ReLU ReLU ReLU
Convolutional Width 1 1 1 1
Dense Activation Function ReLU ReLU ReLU ReLU
Dense Width 512 512 512 512
Normalization None None None None
Discount Factor 0.99 0.99 0.99 0.99
Exploration $\epsilon$ 0.01 0.01 0.01 0.01
Learning Rate 6.25e-5 6.25e-5 0.00005 0.00005
Minimum Replay History 20000 20000 20000 20000
Number of Atoms 0 51 - -
Kappa - - 1.0 1.0
Num tau samples - - 64 64
Num tau prime samples - - 64 64
Num quantile samples - - 32 32
Number of Convolutional Layers 3 3 3 3
Number of Dense Layers 2 2 2 2
Replay Capacity 1000000 1000000 1000000 1000000
Reward Clipping True True True True
Update Horizon 1 3 3 3
Update Period 4 4 4 4
Weight Decay 0 0 0 0
Sticky Actions True True True True
Tau - - 0 0.03
@@
Dueling True True
Double DQN True True
@@ Appendix C Extra results
Figure 16: Normalized performance across 26 Atari games for DrQ($\epsilon$) (left) and DER (right), with the ResNet architecture (Espeholt et al., 2018) and 4 experts. Soft MoE not only remains generally stable with more training, but also attains higher final performance. We report interquantile mean performance with error bars indicating 95% confidence intervals.
Figure 17: Normalized performance across 17 Atari games for CQL+C51. x-axis represents gradient steps; no new data is collected. Left: 10% and Right: 50% uniform replay. We report IQM with 95% stratified bootstrap CIs (Agarwal et al., 2021)
@@ Appendix D Varying Impala filter sizes
When dealing with small models, it’s common to scale them up to enhance performance. This makes the scaling strategy crucial for balancing accuracy and efficiency. For Convolutional Neural Networks (CNNs), traditional scaling methods usually emphasize model depth, width, and input resolution (Ding et al., 2022), as well as the filter. The default filter size is 3x3 for the Impala CNN, and we ran experiments with and without SoftMoE using 4x4 and 6x6 filters to investigate the filter size scaling benefits. In both cases, SoftMoE outperforms the baseline.
Figure 18: Normalized performance across 20 Atari games with the ResNet architecture. SoftMoE achieves the best results in both scenarios; default filter size (3x3) is increased to (4x4) and (6x6).
@@ Appendix E Measuring runtime
Figure 19: Measuring wall-time versus IQM of human-normalized scores in Rainbow over 20 games. Left: ImpalaCNN and Right: CNN network. Each experiment had 3 independent runs, and the confidence intervals show 95% confidence intervals.
@@ Appendix F Experiments with PPO
Figure 20: Left: Evaluating SAC with SoftMoE on 28 MuJoCo environments and Right: Evaluatin PPO on 9 MuJoCo-Brax environments. SoftMoEs seems to provide no gains nor degradation, whereas TopK seems to degrade performance (consistent with paper’s findings). MuJoCo scores are normalized between 0 and 1000, with 5 seeds each; error bars indicate 95% stratified bootstrap confidence intervals. MuJoCo-Brax scores are normalized with respect to Jesson et al. (2023).

### v3

@@ 2.1 Reinforcement Learning
$V^{\pi}(x):=\underset{\pi,\mathcal{P}}{\mathbb{E}}\left[\sum_{t=i}^{\infty}\gamma^{t}\mathcal{R}\left(x_{t},a_{t}\right)\mid x_{0}=x\right].$ The state-action function $Q^{\pi}$ quantifies the value of first taking action $a$ from state $x$, and then following $\pi$ thereafter: $Q^{\pi}(x,a):=\mathcal{R}(x,a)+\gamma\underset{x^{\prime}\sim\mathcal{P}(x,a)}{\mathbb{E}}V^{\pi}(x^{\prime})$. Every MDP admits an optimal policy $\pi^{*}$ in the sense that $V^{\pi^{*}}:=V^{*}\geq V^{\pi}$ uniformly over $\mathcal{X}$ for all policies $\pi$.
@@ 4 Empirical evaluation
Our implementation, included with this submission, is built on the Dopamine library (Castro et al., 2018) [^3] Dopamine code available at https://github.com/google/dopamine., which adds stochasticity (via sticky actions) to the ALE (Machado et al., 2018). We use the recommendations from Agarwal et al. (2021) for statistically-robust performance evaluations, in particular focusing on interquartile mean (IQM). Every experiment was run for 200M environment steps, unless reported otherwise, with $5$ independent seeds, and we report $95\%$ stratified bootstrap confidence intervals. All experiments were run on NVIDIA Tesla P100 GPUs, and each took on average $4$ days to complete.
@@ 4.1 Soft MoE Helps Parameter Scalability
As Figure 1 demonstrates, Soft MoE provides clear performance gains, and these gains increase with the number of experts; for instance in Rainbow, increasing the number of experts from $1$ to $8$ results in a $20\%$ performance improvement. In contrast, the performance of the base architectures declines as we widen its layer; for instance in Rainbow, as we increase the layer multiplier from $1$ to $8$ there is a performance decrease of around $40\%$. This is a finding consistent with prior work demonstrating the difficulty in scaling up deep RL networks (Farebrother et al., 2022; Taiga et al., 2022; Schwarzer et al., 2023). Top1-MoE seems to provide gains when incorporated into Rainbow, but fails to exhibit the parameter-scalability we observe with Soft MoE.
@@ 5.1 Offline RL
We begin by incorporating MoEs in offline RL, where agents are trained on a fixed dataset without environment interactions. Following prior work (Kumar et al., 2021b), we train the agents over $17$ games and $5$ seeds for $200$ iterations, where $1$ iteration corresponds to 62,500 gradient updates. We evaluated on datasets composed of $5\%$, $10\%$, $50\%$ of the samples (drawn randomly) from the set of all environment interactions collected by a DQN agent trained for $200$M steps (Agarwal et al., 2020). In Figure 11 we observe that combining Soft MoE with two modern offline RL algorithms (CQL (Kumar et al., 2020) and CQL+C51 (Kumar et al., 2022)) attains the best aggregate final performance. Top1-MoE provides performance improvements when used in conjunction with CQL+C51, but not with CQL. Similar improvements can be observed when using 10% of the samples (see Figure 17).
@@ 6 Related Work
Another line of research demonstrates that RL networks can be trained with a high sparsity level ($\sim$90%) without loss in performance (Tan et al., 2022; Sokar et al., 2022; Graesser et al., 2022; Ceron et al., 2024). These observations call for techniques to better utilize the network parameters in RL training, such as using MoEs, which we show decreases dormant neurons drastically over multiple tasks and architectures. To enable scaling networks in deep RL, prior works focus on algorithmic methods (Farebrother et al., 2022; Taiga et al., 2022; Schwarzer et al., 2023; Farebrother et al., 2024). In contrast, we focus on alternative network topologies to enable robustness towards scaling.
@@ 7 Discussion and Conclusion
Figure 12: Normalized performance across $26$ Atari games for DrQ($\epsilon$) (left) and DER (right), with the ResNet architecture (Espeholt et al., 2018) and $8$ experts (see Figure 16, for $4$ experts). Soft MoE not only remains generally stable with more training, but also attains higher final performance. We report interquantile mean performance with error bars indicating $95\%$ confidence intervals.
@@ Acknowledgements
The authors would like to thank Gheorghe Comanici, Owen He, Alex Muzio, Adrien Ali Taïga, Rishabh Agarwal, Hugo Larochelle, Ayoub Echchahed, and the rest of the Google DeepMind Montreal team for valuable discussions during the preparation of this work; Gheorghe deserves a special mention for providing us valuable feedback on an early draft of the paper. We thank the anonymous reviewers for their valuable help in improving our manuscript. We would also like to thank the Python community (Van Rossum & Drake Jr, 1995; Oliphant, 2007) for developing tools that enabled this work, including NumPy (Harris et al., 2020), Matplotlib (Hunter, 2007), Jupyter (Kluyver et al., 2016), Pandas (McKinney, 2013) and JAX (Bradbury et al., 2018).
@@ Appendix A Experimental details
Unless otherwise specified, in all experiments below we report the interquantile mean after $40$ million environment steps; error bars indicate $95\%$ stratified bootstrap confidence intervals (Agarwal et al., 2021). Most of our experiments were run with $20$ games from the ALE suite (Bellemare et al., 2013a), as suggested by Fedus et al. (2020). However, for the Atari $100k$ agents (subsection 5.2), we used the standard set of $26$ games (Kaiser et al., 2020) to be consistent with the benchmark. Finally, we also ran some experiments with the full set of $60$ games. The specific games are detailed below.
@@ Appendix B Hyper-parameters list
Default hyper-parameter settings for DER (Van Hasselt et al., 2019) and DrQ($\epsilon$) (Kaiser et al., 2020; Agarwal et al., 2021) agents. Table 1 shows the default values for each hyper-parameter across all the Atari games.
@@
Exploration $\epsilon$ 0.01 0.01
Exploration $\epsilon$ decay 2000 5000
@@
Default hyper-parameter settings for DQN (Mnih et al., 2015) and Rainbow (Hessel et al., 2018) agents. Table 2 shows the default values for each hyper-parameter across all the Atari games.
Table 2: Default hyper-parameters setting for DQN and Rainbow agents.
@@
Hyper-parameter DQN Rainbow
Adam’s ($\epsilon$) 1.5e-4 1.5e-4
Adam’s learning rate 6.25e-5 6.25e-5
Batch Size 32 32
Conv. Activation Function ReLU ReLU
Convolutional Width 1 1
Dense Activation Function ReLU ReLU
Dense Width 512 512
Normalization None None
Discount Factor 0.99 0.99
Exploration $\epsilon$ 0.01 0.01
Exploration $\epsilon$ decay 250000 250000
Minimum Replay History 20000 20000
Number of Atoms 0 51
Number of Convolutional Layers 3 3
Number of Dense Layers 2 2
Replay Capacity 1000000 1000000
Reward Clipping True True
Update Horizon 1 3
Update Period 4 4
Weight Decay 0 0
Sticky Actions True True
@@
Dueling False True
Double DQN False True
@@ Appendix C Extra results
Figure 16: Normalized performance across $26$ Atari games for DrQ($\epsilon$) (left) and DER (right), with the ResNet architecture (Espeholt et al., 2018) and $4$ experts. Soft MoE not only remains generally stable with more training, but also attains higher final performance. We report interquantile mean performance with error bars indicating $95\%$ confidence intervals.
Figure 17: Normalized performance across $17$ Atari games for CQL+C51. x-axis represents gradient steps; no new data is collected. Left: $10\%$ and Right: $50\%$ uniform replay. We report IQM with $95\%$ stratified bootstrap CIs (Agarwal et al., 2021)
@@ Appendix D Varying Impala filter sizes
When dealing with small models, it’s common to scale them up to enhance performance. This makes the scaling strategy crucial for balancing accuracy and efficiency. For Convolutional Neural Networks (CNNs), traditional scaling methods usually emphasize model depth, width, and input resolution (Ding et al., 2022), as well as the filter. The default filter size is $3$x3333 for the Impala CNN, and we ran experiments with and without SoftMoE using $4$x4444 and $6$x6666 filters to investigate the filter size scaling benefits. In both cases, SoftMoE outperforms the baseline.
Figure 18: Normalized performance across $20$ Atari games with the ResNet architecture. SoftMoE achieves the best results in both scenarios; default filter size ($3$x3333) is increased to ($4$x4444) and ($6$x6666).
@@ Appendix E Measuring runtime
Figure 19: Measuring wall-time versus IQM of human-normalized scores in Rainbow over $20$ games. Left: ImpalaCNN and Right: CNN network. Each experiment had $3$ independent runs, and the confidence intervals show $95\%$ confidence intervals.
@@ Appendix F Experiments with PPO
Figure 20: Left: Evaluating SAC with SoftMoE on $28$ MuJoCo environments and Right: Evaluatin PPO on $9$ MuJoCo-Brax environments. SoftMoEs seems to provide no gains nor degradation, whereas TopK seems to degrade performance (consistent with paper’s findings). MuJoCo scores are normalized between 00 and $1000$, with $5$ seeds each; error bars indicate $95\%$ stratified bootstrap confidence intervals. MuJoCo-Brax scores are normalized with respect to Jesson et al. (2023).